An object is released from rest and rolls down an inclined plane, requiring 4 s to cover a distance of 100 cm.
a. What is its acceleration?
b. What is the velocity after the object reaches the 100 cm mark?
1
Expert's answer
2020-10-19T13:22:52-0400

a. Assuming it started moving from rest, the average acceleration is


"a=\frac{2d}{t^2}=\frac{2\cdot1}{4^2}=0.125\text{ m/s}^2."

b. The velocity (assuming the body kept this value of acceleration) is


"v=\sqrt{2ad}=\sqrt{2\cdot0.125\cdot1}=0.5\text{ m/s}^2."